OpenStudy (anonymous):

4u^2/u^2-36 divided by -2u/u^2+8u+12

OpenStudy (anonymous):

((4u^2) * (u^2 + 8u +12)) / ((u^2 -36) * (-2u))

OpenStudy (anonymous):

\[\frac{4u^2}{u^2-36} / \frac{-2u}{u^2+8u+12}\] \[=\frac{4u^2}{(u+6)(u-6)} * \frac{(u+6)(u+2)}{-2u}\] \[=\frac{-2u(u+2)}{(u-6)}\]